OpenStudy (anonymous):

Simplify: 3a^2b^3 / 9ab^4

OpenStudy (anonymous):

a/3b

OpenStudy (anonymous):

\[3a ^{2}b ^{3}\div 9ab ^{4}\] Dont quote me on this but I think the answer \[3ab ^{2}\div 3b^{3}\] I'm really not sure though. Just trying to help. : ^ )

OpenStudy (anonymous):

3a^2b^3 / 9ab^4=3/9*(a^2/a)*(b^3/B^4) =(1/3)*a*(1/b)=a/3b